# Unpacking Approaches to Learning and Teaching Machine Learning in K-12 Education: Transparency, Ethics, and Design Activities

### v1

@@ Abstract.
In this conceptual paper, we review existing literature on artificial intelligence/machine learning (AI/ML) education to identify three approaches to how learning and teaching ML could be conceptualized. One of them, a data-driven approach, emphasizes providing young people with opportunities to create data sets, train, and test models. A second approach, learning algorithm-driven, prioritizes learning about how the learning algorithms or engines behind how ML models work. In addition, we identify efforts within a third approach that integrates the previous two. In our review, we focus on how the approaches: (1) glassbox and blackbox different aspects of ML, (2) build on learner interests and provide opportunities for designing applications, (3) integrate ethics and justice. In the discussion, we address the challenges and opportunities of current approaches and suggest future directions for the design of learning activities.
@@ 1. Introduction
While researchers have been investigating how to introduce young people to Artificial Intelligence/Machine Learning (AI/ML) ideas for decades (Goldstein and Papert, 1977; Kahn, 1977), it is only during the last few years that AI/ML education has gained momentum (Kahn and Winters, 2021). This momentum is a product of the comeback of machine learning methods, this time powered by increasing computing capacity and increasingly large datasets (Touretzky et al., 2019), the establishment of guidelines and principles (Touretzky et al., 2019), and the design of tools that enable novices to create models with small data sets (Carney et al., 2020; Druga, 2018). Today, in light of the popularization of large language models and generative models, young people interact with ML every day; governments call for increasing AI/ML education (U.S. Department of Education, 2023; Commission, 2020), curriculum providers update their offerings (Code.org, 2023), and teachers scramble to integrate AI/ML content in their classes.
@@ 2.1. ML Education in K-12
In their conceptualization of the five big ideas for AI education (see Figure 1), Touretzky et al. (Touretzky et al., 2019) highlight the importance of machine learning. Machine learning enables algorithmic systems to learn behaviors without explicit programming (Tedre et al., 2021). This functionality is achieved by a learning algorithm modifying the internal representations of a reasoning model (e.g., decision tree or neural network), allowing it to learn new behaviors. To effectively narrow the algorithm’s choices, the reasoning model requires a large amount of training data. Once the machine learning algorithm has created the reasoner, it can be used to solve problems and make decisions on new data (Touretzky et al., 2023). As such, machine learning provides numerous opportunities for K-12 students to experiment with training data, learning algorithms, and the design of applications.
@@
Furthermore, as Tedre and colleagues (Tedre et al., 2021) argue, the integration of machine learning into computing education poses challenges to computational thinking practices that have been adopted in the past decade. This is a product of the opacity of ML algorithms, the data driven and data intensive nature of ML. Similarly, Shapiro and Tissenbaum highlight that in contrast to traditional paradigms in computing education, ML is much more empirical, involving learners in conducting experiments, formulating hypotheses, and evaluating models based on predictions (Shapiro and Tissenbaum, 2019). As such, designing learning activities requires making important decisions on how to address the opacity of ML systems, how to engage students in training and evaluating models, and finding accessible entry points.
@@ 2.2.1. Designing Applications
As Resnick and colleagues (Resnick et al., 1999) argue, tools that glassbox the inner workings of computing can support learners in designing applications that would otherwise be inaccessible. Designing applications in computing education has a long history that goes back to studies on LOGO (Harel, 1991; Kafai, 1995). This approach of having K-12 students create projects that relate to their interests, instead of completing pre-determined problem sets, has gained traction over the past decade (Waite and Sentance, 2021; Oleson et al., 2020). In ML education, for instance, learning activities may blackbox and glassbox different aspects of the ML pipeline to support students to move beyond just training models in solation to create models to be used in fully developed applications.
@@ 2.2.2. Algorithmic Justice and Ethics
More recently, Dixon, Hsi and Oh (Dixon et al., 2022) revisited the ideas of providing black and glassbox scaffoldings to propose that transparency should not be solely focused on technical disciplinary knowledge and practices; instead, it should also consider the histories, externalities, and possible futures of computing technologies. Here it is important for youth to engage with ideas of algorithmic justice. Algorithmic justice considers that ML systems have implications on individuals and communities that could perpetuate harm in unjust ways, disproportionately impacting vulnerable populations (Birhane, 2021). As such, making ML blackboxes transparent for learners involves providing scaffolds to think through how functional and critical issues are intertwined, consider the possible ways in which ML applications could be used, and take into account their ethical, social, and environmental implications.
@@ 4.1. Data-driven approach
The most common approach to AI/ML education in K-12 is data-driven, having students build datasets to train models. This approach glassboxes how data shapes model performance and blackboxes the role of learning algorithms in the ML pipeline. We reviewed 40 different studies that take a data-centric approach. Most studies within the past five years have focused on having young people create and label datasets for classification tasks (e.g., (Ng et al., 2022; Jordan et al., 2021; Druga et al., 2022a; Arastoopour Irgens et al., 2022; Hjorth, 2021). This commonly has young people engage with what Fiebrink (Fiebrink, 2019) calls the Interactive Machine Learning approach which involves iteratively ideating a system and its desired behavior, implementing it, observing its behavior, and comparing the observed behavior with the desired behaviors. Here users create small data sets that can be easily and quickly refined and edited to retrain models and improve their performance (Zimmermann-Niefield et al., 2019). However, Hitron et al. (Hitron et al., 2019) show that to build basic conceptual understanding of machine learning through a data-driven approach, solely labeling data or just evaluating models is not enough. In a study with children they found that engaging with the full pipeline of creating a dataset, training and testing a model can lead to significant improvements in conceptual understanding when compared with just labeling data or evaluating outcomes. More recently, researchers have articulated data design practices for novices building ML models based on expert practices (Tseng et al., 2024). These include 1) incorporating dataset diversity, 2) evaluating model performance and its relationship to data, 3) balancing datasets, and 4) inspecting for data quality. The data-driven approach promotes what Vartianen and colleagues (Vartiainen et al., 2021) call “data-driven reasoning and design” which involves thinking about decisions made in the design of datasets to explain the behavior of machine learning systems. As such in this approach model behaviors are ascribed to the design of datasets.
@@
This approach has been used in diverse contexts: in and out of school, with very young children (Vartiainen et al., 2020a; Williams et al., 2019b) and adolescents (e.g., (Tseng et al., 2024; Sabuncuoglu, 2020)), in computing classes and workshops (Arastoopour Irgens et al., 2022) and in other disciplines such as social studies (Hjorth, 2021) or chemestry (Martin et al., 2024). For example Lin and colleagues (Lin et al., 2020) involved young children in designing a conversational agent by providing data related to specific animals. Sabuncouglu and colleagues (Sabuncuoglu, 2020), on the other hand, designed and implemented a year-long curriculum to introduce adolescents to machine learning, in which they designed data sets for rock, paper, scissors classifiers, and a personal project to address environmental issues. One study goes beyond classification tasks by having children provide data (melodies) for a generative music model (Williams et al., 2019b). Other studies build on unplugged computing activities, using a card game for young people to reflect on the context of data, data inputs and data ethics (Bilstrup et al., 2020).
@@ 4.1.1. Tools & Curricula
Studies that approach ML education from a data-driven perspective introduced new tools to support learners in building models and presented curricular interventions.
Twenty six studies introduced new tools that enable learners to build and create models. These go from extensions of Scratch to support face, body, and hand recognition and the possibility of including Teachable Machine models (Jordan et al., 2021; Rodríguez-García et al., 2021) to novel tools for using natural language processing in social studies classrooms (Hjorth, 2021; Jiang et al., 2022), to gesture classification tools. One tool involves a physical tangible artifact that can be used by children to create datasets using drawings, train models and test them (Kaspersen et al., 2021). Hjorth (Hjorth, 2021) presented Text Tagging, an application for learners to use NLP techniques on social media posts in their social studies classes. Gesture classification tools include Gest (Hitron et al., 2018), PlushPal (Tseng et al., 2021) and AlpacaML (Zimmermann-Niefield et al., 2019) which all support learners in building and testing ML models that use accelerometer sensor data. Despite their similarities, each of these tools supports different kinds of engagement with gesture classification from engaging young children in telling stories about their plushies (Tseng et al., 2021) to teenagers in creating sports related projects (Zimmermann-Niefield et al., 2019). More advanced tools, such as StoryQ (Jiang et al., 2022), also support learners to visualize and explore patterns in data. Unplugged tools included a card game about data for ML models with cards that determine the context of data, data inputs, and prompts about how data is collected and its ethics (Bilstrup et al., 2020).
Nineteen studies presented curricular activities to engage students with data-driven ML topics in distinct ways. Ng et al. (Ng et al., 2022) for example present a curriculum on storytelling and ML and Hjorth (Hjorth, 2021) introduce ML activities for social studies. Other curricula take a problem-based approach to have youth address socio-technical problems using ML (Krakowski et al., 2022). Some efforts include curricular activities for at-home learning with families (Druga et al., 2022a; Tseng et al., 2023). Other scaffolds for learners included activity cards, sample projects and worksheets (Jordan et al., 2021). While some efforts are short after school activities (Arastoopour Irgens et al., 2022) others involve year long curricula (Rodríguez-García et al., 2021).
@@ 4.1.2. Making models in the context of designing applications
Of the 40 studies we reviewed, 19 studies had students train models in the context of building applications and 24 centered on building models in isolation. Building applications in this context involves creating datasets and training models and using those models in applications usually designed by learners themselves. As such, learners have to engage both with machine learning as well as traditional computing education practices and concepts (see (Tedre et al., 2021) for a discussion of the difference between traditional computing and ML education).
Sixteen of these studies involved students in model building in the context of creating personally relevant projects. Poseblocks (Jordan et al., 2021) for example, involved students in creating AR filters related to their personal interests Several studies have k-12 students build models to use in personally relevant Scratch projects (Zimmermann-Niefield et al., 2020; Ng et al., 2022; Jordan et al., 2021). Examples here include children making a project to help their siblings learn their ABCs (Jordan et al., 2021). However, as Zimmermann-Niefield and colleagues (Zimmermann-Niefield et al., 2020) note sometimes the projects and models can be incoherent with, in the case of their study, gestures being classified that had little relationship to narratives and characters of Scratch projects.
Other studies involved learners in building data sets, training and testing models and brainstorming applications in which these could be used (Druga et al., 2022a; Toivonen et al., 2020; Vartiainen et al., 2020a, 2021). Here learners created paper prototypes of applications that could use their models. In a few of these studies the prototypes were then developed by software engineers to have learners test them (Vartiainen et al., 2020b, 2021). While this is not scalable and applicable in most k-12 classrooms, it provided a rich opportunity for students to see how the models they trained could be used in real applications.
@@ 4.1.3. Testing
Most studies focus on having learners prepare training datasets for models with little attention paid to creating testing datasets and testing. Indeed, most testing efforts presented in current research involve live-testing, that is using a trained model to classify new inputs in real time (Toivonen et al., 2020; Vartiainen et al., 2021; Druga and Ko, 2021). Studies show that live-testing can be beneficial for youth to take perspective on their models, build hypotheses, and iteratively improve model performance (Morales-Navarro et al., 2024b; Vartiainen et al., 2021; Druga and Ko, 2021); yet testing activities seem to work better when scaffolded by the learning activities (Zimmermann-Niefield et al., 2020). Engaging with testing data sets is only present in a few studies. In Popbots (Williams et al., 2019b), for example, young children experimented with different training datasets to see their impact on accuracy in classifying a testing set. In studies by Tseng et al. (Tseng et al., 2023, 2024) youth created testing data sets and determined the accuracy of their models by class or label. Other studies mention evaluation more broadly. Krakowski et al. (Krakowski et al., 2022), for instance, frame evaluation as involving the following: determining if ML is appropriate for the task at hand, reviewing and questioning the data used to train a model, considering the affordances and limitations of the design of ML systems, and taking into account the foreseeable impact of ML decisions.
@@ 4.1.4. Algorithmic justice and ethics
Few studies that take on a data-driven approach (only 12 out of 40) addressed issues of algorithmic justice. At the same time, while the studies acknowledged the importance of learning about ethics and reducing algorithmic harms, these topics were often presented as something that should be learned separate from skills and concepts. Here, ethics are discussed in relation to commercial applications instead of being applied to learner-designed applications. For example, some studies use ethical matrices for children to redesign YouTube (Sabuncuoglu, 2020) or foster discussions of ethics while redesigning voice assistants (Druga et al., 2022a). A common activity several studies involves watching videos of Joy Buolamwini (Jordan et al., 2021; Arastoopour Irgens et al., 2022)).
Some approaches to issues of justice may be problematic. One study, for example, showed that some children’s concerns may align with long termist sci-fi-inspired ideas (Ng et al., 2022). This shows that it is important to scaffold conversations about justice and ethics to ensure learners can have informed discussions. Another issue that may be problematic is the idea of unbiased ML systems, which a teacher in one study voiced (Ng et al., 2022). Some studies foster AI4SocialGood ideas (e.g., (Krakowski et al., 2022; Arastoopour Irgens et al., 2022), here it is important to avoid falling into the trap of technosolutionism and to ask if ML is necessary to address the problems at hand.

### v2

@@ Abstract.
In this conceptual paper, we review existing literature on artificial intelligence/machine learning (AI/ML) education to identify three approaches to how learning and teaching ML could be conceptualized. One of them, a data-driven approach, emphasizes providing young people with opportunities to create data sets, train, and test models. A second approach, learning algorithm-driven, prioritizes learning about learning algorithms. In addition, we identify efforts within a third approach that integrates the previous two. In our review, we focus on unpacking how the approaches: (1) glassbox and blackbox different aspects of ML, (2) build on learner interests and provide opportunities for designing applications, (3) integrate ethics and justice. In the discussion, we address the challenges and opportunities of current approaches and suggest future directions for the design of tools and learning activities.
@@ 1. Introduction
While researchers have been investigating how to introduce young people to Artificial Intelligence/Machine Learning (AI/ML) ideas for decades (Goldstein and Papert, 1977; Kahn, 1977), it is only during the last few years that AI/ML education has gained momentum (Kahn and Winters, 2021). This momentum has been a product of the comeback of machine learning methods, this time powered by increasing computing capacity and increasingly large datasets (Touretzky et al., 2019), the establishment of guidelines and principles (Touretzky et al., 2019), and the design of tools that enable novices to create models with small data sets (Carney et al., 2020; Druga, 2018). Today, in light of the popularization of large language models and generative models, young people interact with ML every day; governments call for increasing AI/ML education (U.S. Department of Education, 2023; Commission, 2020), curriculum providers update their offerings (Code.org, 2023), and teachers scramble to integrate AI/ML content in their classes.
@@ 2.1. ML Education in K-12
In their conceptualization of the five big ideas for AI education (see Figure 1), Touretzky et al. (Touretzky et al., 2019) highlight the role of machine learning methods as the driver of AI adoption over the past fifteen years. Machine learning enables algorithmic systems to learn behaviors without explicit programming (Tedre et al., 2021). This functionality is achieved by a learning algorithm modifying the internal representations of a reasoning model (e.g., decision tree or neural network), allowing it to learn new behaviors. To effectively narrow the algorithm’s choices, the reasoning model requires a large amount of training data. Once the machine learning algorithm has created the reasoner, it can be used to solve problems and make decisions on new data (Touretzky et al., 2023). As such, machine learning provides numerous opportunities for K-12 students to experiment with training data, learning algorithms, and the design of applications.
@@
Furthermore, as Tedre and colleagues (Tedre et al., 2021) argue, the integration of machine learning into computing education poses challenges to computational thinking practices that have been adopted in the past fifteen years. This is a product of the opacity of ML algorithms, the data driven and data intensive nature of ML. Similarly, Shapiro and Tissenbaum highlight that in contrast to traditional paradigms in computing education, ML is much more empirical, involving learners in conducting experiments, formulating hypotheses, and evaluating models based on predictions (Shapiro and Tissenbaum, 2019). As such, designing learning activities requires making important decisions on how to address the opacity of ML systems, how to engage students in training and evaluating models, and finding accessible entry points.
@@ 2.2.1. Designing Applications
As Resnick and colleagues (Resnick et al., 1999) argue, tools that glassbox the inner workings of computing can support learners in designing applications that would otherwise be inaccessible. Designing applications in computing education has a long history that goes back to studies on LOGO (Harel, 1991; Kafai, 1995). This approach of having K-12 students create projects that relate to their interests, instead of completing pre-determined problem sets, has gained traction over the past decade (Waite and Sentance, 2021; Oleson et al., 2020). In ML education, for instance, learning activities may blackbox and glassbox different aspects of the ML pipeline to support students to move beyond just training models in isolation to create models to be used in fully developed applications.
@@ 2.2.2. Algorithmic Justice and Ethics
More recently, Dixon, Hsi and Oh (Dixon et al., 2022) revisited the ideas of providing blackbox and glassbox scaffoldings to propose that transparency should not be solely focused on technical disciplinary knowledge and practices; instead, it should also consider the histories, externalities, and possible futures of computing technologies. Here it is important for youth to engage with ideas of algorithmic justice. Algorithmic justice considers that ML systems have implications on individuals and communities that could perpetuate harm in unjust ways, disproportionately impacting vulnerable populations (Birhane, 2021). As such, making ML blackboxes transparent for learners involves providing scaffolds to think through how functional and critical issues are intertwined, consider the possible ways in which ML applications could be used, and take into account their ethical, social, and environmental implications.
@@ 4.1. Data-driven approach
The most common approach to AI/ML education in K-12 is data-driven, having students build datasets to train models. This approach glassboxes how data shapes model performance and blackboxes the role of learning algorithms in the ML pipeline. We reviewed 40 different studies that took a data-centric approach. Most studies within the past five years have focused on having young people create and label datasets for classification tasks (e.g., (Ng et al., 2022; Jordan et al., 2021; Druga et al., 2022a; Arastoopour Irgens et al., 2022; Hjorth, 2021). This approach engages young people with what Fiebrink (Fiebrink, 2019) calls ”Interactive Machine Learning” which involves iteratively ideating a system and its desired behavior, implementing it, observing its behavior, and comparing the observed behavior with the desired behaviors. Here users create small data sets that can be easily and quickly refined and edited to retrain models and improve their performance (Zimmermann-Niefield et al., 2019). However, Hitron et al. (Hitron et al., 2019) show that to build basic conceptual understanding of machine learning through a data-driven approach, solely labeling data or just evaluating models is not enough. In a study with children they found that engaging with the full pipeline of creating a dataset, training and testing a model can lead to significant improvements in conceptual understanding when compared with just labeling data or evaluating outcomes. More recently, researchers have articulated data design practices for novices building ML models based on expert practices (Tseng et al., 2024). These include 1) incorporating dataset diversity, 2) evaluating model performance and its relationship to data, 3) balancing datasets, and 4) inspecting for data quality. The data-driven approach promotes what Vartianen and colleagues (Vartiainen et al., 2021) call “data-driven reasoning and design” which involves thinking about decisions made in the design of datasets to explain the behavior of machine learning systems. As such in this approach model behaviors are ascribed to the design of datasets.
@@
This approach has been used in diverse contexts: in and out of school, with very young children (Vartiainen et al., 2020a; Williams et al., 2019b) and adolescents (e.g., (Tseng et al., 2024; Sabuncuoglu, 2020)), in computing classes and workshops (Arastoopour Irgens et al., 2022) and in other disciplines such as social studies (Hjorth, 2021) or chemistry (Martin et al., 2024). For example Lin and colleagues (Lin et al., 2020) involved young children in designing a conversational agent by providing data related to specific animals. Sabuncouglu and colleagues (Sabuncuoglu, 2020), on the other hand, designed and implemented a year-long curriculum to introduce adolescents to machine learning, in which they designed data sets for rock, paper, scissors classifiers, and a personal project to address environmental issues. One study went beyond classification tasks by having children provide data (melodies) for a generative music model (Williams et al., 2019b). Other studies built on unplugged computing activities, using a card game for young people to reflect on the context of data, data inputs and data ethics (Bilstrup et al., 2020).
@@ 4.1.1. Tools & Curricula
Studies that approached ML education from a data-driven perspective introduced tools to support learners in building models and presented curricular interventions.
Twenty six studies introduced new tools that enable learners to build and create models. These included extensions of Scratch to support face, body, and hand recognition and the possibility of including Teachable Machine models (Jordan et al., 2021; Rodríguez-García et al., 2021), novel tools for using natural language processing in social studies classrooms (Hjorth, 2021; Jiang et al., 2022), and gesture classification tools. One tool involved a physical tangible artifact that could be used by children to create datasets using drawings, train models and test them (Kaspersen et al., 2021). Hjorth (Hjorth, 2021) presented Text Tagging, an application for learners to use NLP techniques on social media posts in their social studies classes. Gesture classification tools included Gest (Hitron et al., 2018), PlushPal (Tseng et al., 2021) and AlpacaML (Zimmermann-Niefield et al., 2019) which all supported learners in building and testing ML models that use accelerometer sensor data. Despite their similarities, each of these tools supports different kinds of engagement with gesture classification from supporting young children in telling stories about their plushies (Tseng et al., 2021) to teenagers in creating sports related projects (Zimmermann-Niefield et al., 2019). More advanced tools, such as StoryQ (Jiang et al., 2022), also support learners to visualize and explore patterns in data. Unplugged tools included a card game about data for ML models, with cards that determined the context of data, data inputs, and prompts about how data is collected and its ethics (Bilstrup et al., 2020).
Nineteen studies presented curricular activities to engage students with data-driven ML topics in distinct ways. Ng et al. (Ng et al., 2022) for example presented a curriculum on storytelling and ML and Hjorth (Hjorth, 2021) introduced ML activities for social studies. Other curricula took a problem-based approach to have youth address socio-technical problems using ML (Krakowski et al., 2022). Some efforts included curricular activities for at-home learning with families (Druga et al., 2022a; Tseng et al., 2023). Scaffolds for learners included activity cards, sample projects and worksheets (Jordan et al., 2021). While some efforts were short after school activities (Arastoopour Irgens et al., 2022) others involved year long curricula (Rodríguez-García et al., 2021).
@@ 4.1.2. Making models in the context of designing applications
Of the 40 studies we reviewed, 19 studies had students train models in the context of building applications and 24 centered on building models in isolation. Building applications in this context involved creating datasets and training models and using those models in applications usually designed by learners themselves. As such, learners had to engage both with machine learning as well as traditional computing education practices and concepts (see (Tedre et al., 2021) for a discussion of the difference between traditional computing and ML education).
Sixteen of these studies involved students in model building in the context of creating personally relevant projects. Poseblocks (Jordan et al., 2021) for example, involved students in creating AR filters related to their personal interests. Several studies had students build models to use in personally relevant Scratch projects (Zimmermann-Niefield et al., 2020; Ng et al., 2022; Jordan et al., 2021). Examples here included children making a project to help their siblings learn their ABCs (Jordan et al., 2021). However, as Zimmermann-Niefield and colleagues (Zimmermann-Niefield et al., 2020) noted sometimes the projects and models can be incoherent, in the case of their study, with classified gestures having little relationship to narratives and characters of Scratch projects.
Other studies involved learners in building data sets, training and testing models and brainstorming applications in which these could be used (Druga et al., 2022a; Toivonen et al., 2020; Vartiainen et al., 2020a, 2021). Here learners created paper prototypes of applications that could use their models. In a couple of these studies the prototypes were then developed by software engineers to have learners test them (Vartiainen et al., 2020b, 2021). While this is not scalable and applicable in most k-12 classrooms, it provided a rich opportunity for students to see how the models they trained could be used in real applications.
@@ 4.1.3. Testing
Most studies focused on having learners prepare training datasets for models with little attention paid to creating testing datasets and testing. Indeed, most testing efforts presented in current research involved live-testing, that is using a trained model to classify new inputs in real time (Toivonen et al., 2020; Vartiainen et al., 2021; Druga and Ko, 2021). Studies show that live-testing can be beneficial for youth to take perspective on their models, build hypotheses, and iteratively improve model performance (Morales-Navarro et al., 2024b; Vartiainen et al., 2021; Druga and Ko, 2021); yet testing activities seem to work better when explicitly scaffolded by instructors (Zimmermann-Niefield et al., 2020). Engaging with testing data sets was only present in a few studies. In Popbots (Williams et al., 2019b), for example, young children experimented with different training datasets to see their impact on accuracy in classifying a testing set. In studies by Tseng et al. (Tseng et al., 2023, 2024) youth created testing data sets and determined the accuracy of their models by class or label. Other studies mention evaluation more broadly. Krakowski et al. (Krakowski et al., 2022), for instance, framed evaluation as a process that involves: determining if ML is appropriate for the task at hand, reviewing and questioning the data used to train a model, considering the affordances and limitations of the design of ML systems, and taking into account the foreseeable impact of ML decisions.
@@ 4.1.4. Algorithmic justice and ethics
Few studies that took on a data-driven approach (only 12 out of 40) addressed issues of algorithmic justice. At the same time, while the studies acknowledged the importance of learning about ethics and reducing algorithmic harms, these topics were often presented as something that should be learned separate from skills and concepts. Here, ethics were discussed in relation to commercial applications instead of being applied to learner-designed applications. For example, some studies used ethical matrices for children to redesign YouTube (Sabuncuoglu, 2020) or fostered discussions of ethics while redesigning voice assistants (Druga et al., 2022a). A common activity in several studies involved watching videos of Joy Buolamwini (Jordan et al., 2021; Arastoopour Irgens et al., 2022)).
Some approaches to issues of justice may be problematic. One study, for example, showed that some children’s concerns may align with long termist sci-fi-inspired ideas (Ng et al., 2022). This shows that it is important to scaffold conversations about justice and ethics to ensure learners can have informed discussions. Another issue that may be problematic is the idea of unbiased ML systems, which a teacher in one study voiced (Ng et al., 2022). Some studies foster AI4SocialGood ideas (e.g., (Krakowski et al., 2022; Arastoopour Irgens et al., 2022)), here it is important to avoid falling into the trap of technosolutionism and to ask if ML is necessary to address the problems at hand.